[Brahms-l] Shift report 20050613 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Jun 13 2005 - 07:58:37 EDT
Shifter leader was OG
Shifters were 


1.Beam was ~4.5h.
2. 2 PED and 10 physics runs. There are shorts runs due to RunErrors(The spool output streem is missing)
3. No visible problems with detector.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jun 13 07:58:53 2005

This archive was generated by hypermail 2.1.8 : Mon Jun 13 2005 - 07:59:05 EDT